a grocer mixed nuts worth $4.25 per pound with nuts worth $2.75 per pound. how many pounds of each kind did he use to make a mixture of 60 pounds to sell at $3,75 per pound







Found 2 solutions by Edwin McCravy, addingup:
Answer by Edwin McCravy(20056)


                     Pounds   Price    Dollar
                       of      per      value
                      nuts    pound   of nuts
    Expensive nuts     x       4.25     4.25x 
        Cheap nuts     y       2.75     2.75y
---------------------------------------------
Medium-priced nuts    60       3.75     3.75(60)

The equations are



Simplify and solve.  If you have trouble, you can
tell me in the thank-you note form below, and I'll
get back to you by email.  No charge ever.  I'm an
old retired math prof and I just do this for fun. :)

Edwin

Answer by addingup(3677)

Let's call the nuts at 4.25 x and those at 2.75 y
x+y = 60, if we subtract x from both sides:
y = 60-x
Next:
4.25x+2.75y = 3.75(60) substitute for y:
4.25x+2.75(60-x) = 3.75(60)
4.25x+165-2.75x = 225
1.5x = 60
x = 40
You need 40 pounds of the nuts at 4.25 and (60-40=20) 20 pounds of the nuts at 2.75